Assume that the market demand and supply curves for soybeans grown in Canada can be represented via the following:

QD = 40?0.5PS QS = 2.5+2.5PS (1) where PS is the soybean price ($/bushel) and QS is the quantity of soybeans produced (de-

nominated in 100 million bushel units).


(a) What is the equilibrium price, PS?, and quantity, Q?S, of soybeans?

(b) Using Excel or some other spreadsheet program verify your answer to (a) graph- ically.
Answer #1

It is given that quantity demanded is QD = 40-0.5PS and quantity supplied is Qs=2.5+2.5PS.

At equilibrium, QD = Qs

Or, 40-0.5PS = 2.5+2.5PS.

Or 3PS = 37.5

Or PS = $12.5

Therefore Quantity demanded = Quantity Supplied = 33.75 (in 100 Million Bushel Units)

This has been done in Excel and the steps are as follows:

  • Fill the price column taking intervals of 2.5 because it’s easier to plot 12.5 as 12.5 is divisible by 2.5
  • Fill out the demand column by clicking on the first QD cell, writing the equation “=40-0.5*(corresponding price cell) [Here my first price cell was A10 and I filled this formula in B10. Once I got the result I dragged it through to B17 and the cells automatically filled up.]
  • Similarly, fill out the supply column by clicking on the first QS cell, writing the equation “=2.5+2.5* (corresponding price cell) [Here my first price cell was A10 and I filled this formula in C10. Once I got the result I dragged it through to C17 and the cells automatically filled up.]
  • Select the cells from A10 to B17 (covering Price and Demand) >>Then click on insert >> Select “scatter” >> select the one with straight lines and rectangular data labels on it.
  • Right click on the graphà Select Data >> Under Legend Entries select “Remove”>> select “Add”>>Give the series name as QD>>select X values by clicking on the demand values (B10 through B17)>>Select Y values by clicking on PRICE (A10 through A17).
  • Repeat the process from “Add">> Give the series name as QS>> select X values by clicking on the supply values (C10 through C17)>> Select Y values by clicking on PRICE (A10 through A17)>> Click on ok.
  • After your graph is done, right click on the intersection of QD and Qs>> select “add data labels”.
  • Finish off by adding titles to your graph. Left click on the graph and Select “Layout” beside the tab of ‘design’ >> Choose “Chart Title”>>above chart>> enter the chart title in the dialogue box. In my case it’s “Equilibrium Price and Quantity for Soybeans in USA”
  • Select “Axis Titles” right beside “Chart Title” and label the vertical and horizontal axes in the same way using the dialogue boxes. In this case, the X and Y axes are Quantity (in 100 Million Bushels) and Price ($ per Bushel) respectively.
  • Your graph is done!

Equilibrium Price and Quantity for Soybeans in USA Price emand Supply 14 12 15 15 16417.5 17 20 uantity (20o Million Bushel U